40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5258     Entry time: Wed Aug 17 20:14:49 2011
Author: jamie, kiwamu, suresh, jenne, keiko, anamaria 
Type: Update 
Category: General 
Subject: in-vacuum work status, prep for pump 

This afternoon's work:

  • OSEMs were adjusted on all suspended optics.
  • X and Y-arms were aligned to green.
  • Once that was done, the input pointing was adjusted with the PZTs to get the IR beam centered on ITMY and ETMY.
  • Once the input pointing was ok we extracted IPANG and IPPOS.
  • BS, ITMY, PRM, SRM optical levers (oplevs) were extracted.
  • Prepared rails and stops for TMs for morning drag wiping.

TODO before drag wiping:

  • Check full IFO alignment.
  • Readjust OSEMs if needed.
  • Extract ITMX oplev.
ELOG V3.1.3-